OpenStudy (anonymous):

Need Help solve \(\large x'=\begin{bmatrix} -2 & 1\\ 1 &-2 \end{bmatrix}x+\begin{bmatrix} 1\\ 3 \end{bmatrix}e^{-2t}\)

OpenStudy (michele_laino):

I call your function x(t) as below: \[\left(\begin{matrix}a(t) \\ b(t)\end{matrix}\right)\] so I can write these equations: \[a'=-2a+b+2e ^{-2t},b'=a-2b+3e ^{-2t}\]

OpenStudy (anonymous):

aha k i dont mind :)

OpenStudy (michele_laino):

adding to the doubled first equation the second equation, we can write:\[2a'+b'=-3a+5e ^{-2t}\] deriving one time that equation we have: \[2a''+b''=-3a'-10e ^{-2t}\]

OpenStudy (michele_laino):

so: \[b''(t)=-3a'-10e ^{-2t}-2a''\]

OpenStudy (michele_laino):

now I derive the firs equation two times with respect to t, and I can get this: \[a'''=-2a''+b''+4e _{^{-2t}}\] inserting the expression for b'' above, we can write: \[a'''+4a''+3a'=-10e ^{-2t}\] which can be easily soved

OpenStudy (anonymous):

i see ok so by solving characteristic equation r^3+4r^2+3r=0 r(r^2+4r+3)=0 r(r+1)(r+3)=0 r=0,-1,-3 i would have homogeneous solution \(a_H=c_1+c_2e^{-t}+c_2e^{-3t}\)

OpenStudy (anonymous):

now i have to look for particular solution ?

OpenStudy (michele_laino):

please try with this function: \[a _{p}(t)=A*e ^{-2t}\] where A is a real constant

OpenStudy (michele_laino):

you should get this: \[a _{p}=-5e ^{-2t}\]

OpenStudy (anonymous):

@Michele_Laino shouldn't that equation in \(a\) be \[a'''+4a''+3a'=-\color{red}6e ^{-2t}~~?\]

OpenStudy (michele_laino):

so, a(t) is: \[a(t)=c _{1}+c _{2}e ^{-t}+c _{3}e ^{-3t}-5e ^{-2t}\]

OpenStudy (michele_laino):

I will check, please wait... @SithsAndGiggles

OpenStudy (michele_laino):

you are right! @SithsAndGiggles it is -6e^(-2t)

OpenStudy (michele_laino):

@Marki please note the post of @SithsAndGiggles on my error of computation

OpenStudy (anonymous):

Fortunately the homogeneous solution doesn't change :)

OpenStudy (michele_laino):

that's right! @SithsAndGiggles

OpenStudy (michele_laino):

so our solution for a(t) should be this: \[a(t)=c _{1}+c _{2}e ^{-t}+c _{3}e ^{-3t}-3e ^{-2t}\]

OpenStudy (anonymous):

@Marki note that elimination is not the only way to obtain a solution. You can also try using undetermined coefficients and the usual eigenvector approach for the homogeneous solution. (I prefer this method because you get \(x_1=a\) and \(x_2=b\) right away.)

OpenStudy (anonymous):

hmm i would assume \(a=Ae^{-2t}\) \(a'=-2Ae^{-2t}\) \(a''=4Ae^{-2t}\) \(a'''=-8Ae^{-2t}\) so -8A+16A-6A=-6 A=-1/3

OpenStudy (michele_laino):

please, I got: -14+16=+2, so A=-3 I think

OpenStudy (michele_laino):

whereas our solution for b(t) is: \[b(t)=-3c _{1}t+c _{2}e ^{-t}-c _{3}e ^{-3t}-e ^{-2t}\]

OpenStudy (anonymous):

this my way to solve , but i was looking for other comfortable methods A is Diagonalizable eigen values and corresponding vectors -1,(1,1) -3,(1,-1) and i get some direct solution with some x(t)=Ty(t)=|1 1 ;1 -1||y1;y2| at the end i got some long solution xD i'll post if u interested

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

i got this (in other method) \(\large a(t)=-2e^{-t} (e^{-t}-1) +c_1e^{-t}-e^{-3t}(e^t-1)+c_2e^{-3t}\) \(\large b(t)=-2e^{-t} (e^{-t}-1) +c_1e^{-t}+e^{-3t}(e^t-1)-c_2e^{-3t}\)

OpenStudy (anonymous):

wew! but i completely got ur method hmm but in exam mostly i'll use this since they might give 3*3

OpenStudy (anonymous):

thanks both @Michele_Laino @SithsAndGiggles

OpenStudy (michele_laino):

thanks! for your matrix method! @Marki

OpenStudy (michele_laino):

thanks! for your correction @SithsAndGiggles

OpenStudy (anonymous):

Perhaps I made a mistake myself, but I'm getting a different solution than what elimination gives. \[{\bf x}'=\begin{pmatrix}-2&1\\1&-2\end{pmatrix}{\bf x}+\begin{pmatrix}1\\3\end{pmatrix}e^{-2t}\] The coefficient matrix has the same eigenvalues/vectors, \(\lambda_1=-3\) with \(\vec{\eta}_1=\begin{pmatrix}-1\\1\end{pmatrix}\) and \(\lambda_2=-1\) with \(\vec{\eta}_2=\begin{pmatrix}1\\1\end{pmatrix}\) (which agrees with Marki's work). The homogeneous solution is then \[{\bf x}_h=C_1\begin{pmatrix}-1\\1\end{pmatrix}e^{-3t}+C_2\begin{pmatrix}1\\1\end{pmatrix}e^{-t}\] For the non-homogeneous part, we set \({\bf x}=Ae^{-2t}\), so that \({\bf x}'=-2Ae^{-2t}\), then we substitute into the original equation: \[\begin{align*}-2Ae^{-2t}&=\begin{pmatrix}-2&1\\1&-2\end{pmatrix}Ae^{-2t}+\begin{pmatrix}1\\3\end{pmatrix}e^{-2t}\\\\ \vec{0}&=\left(\left[\begin{pmatrix}-2&1\\1&-2\end{pmatrix}-2I\right]A+\begin{pmatrix}1\\3\end{pmatrix}\right)e^{-2t}\\\\ \begin{pmatrix}-4&1\\1&-4\end{pmatrix}A&=\begin{pmatrix}-1\\-3\end{pmatrix}\\\\ A&=\frac{7}{15}\begin{pmatrix}1\\1\end{pmatrix} \end{align*}\] So the final solution would be (according to this work) \[{\bf x}=C_1\begin{pmatrix}-1\\1\end{pmatrix}e^{-3t}+C_2\begin{pmatrix}1\\1\end{pmatrix}e^{-t}+\frac{7}{15}\begin{pmatrix}1\\1\end{pmatrix}e^{-2t}\]

OpenStudy (anonymous):

Oh I think I see the mistake, I did \(\color{red}{-2IA}\) instead of \(\color{red}{+2IA}\)...

OpenStudy (anonymous):

It should be \[A=-\begin{pmatrix}3\\1\end{pmatrix}\]

OpenStudy (anonymous):

i thinkits ok method are much imp than solution

OpenStudy (anonymous):

to me or u ?

OpenStudy (anonymous):

Both! It's nice knowing you have the right solution